# Centralization vs. decentralization in multi-robot sweep coverage with ground robots and UAVs

### v2

@@ Abstract
In swarm robotics, decentralized control is often proposed as a more scalable and fault-tolerant alternative to centralized control. However, centralized behaviors are often faster and more efficient than their decentralized counterparts. In any given application, the goals and constraints of the task being solved should guide the choice to use centralized control, decentralized control, or a combination of the two. Currently, the exact trade-offs that exist between centralization and decentralization are not well defined. In this paper, we study comparative performance assessment between centralization and decentralization in the example task of sweep coverage, across five different types of multi-robot control structures: random walk, decentralized with beacons, hybrid formation control using self-organizing hierarchy, centralized formation control, and predetermined. In all five approaches, the coverage task is completed by a group of ground robots. In each approach, except for the random walk, the ground robots are assisted by UAVs, acting as supervisors or beacons. We compare the approaches in terms of three performance metrics for which centralized approaches are expected to have an advantage—coverage completeness, coverage uniformity, and sweep completion time—and two metrics for which decentralized approaches are expected to have an advantage—scalability (4, 8, or 16 ground robots) and fault tolerance (0%, 25%, 50%, or 75% ground robot failure). Finally, we discuss future work on combining the advantages of both in one system.
@@ 1 Introduction
When developing multi-robot systems, a fundamental design choice is whether to use centralization, decentralization, or some combination of the two. Ideally, one would be able to guarantee that a selected control approach is the top performer for any targeted objectives. However, current methods to assess performance do not span different system types, making direct comparisons difficult. For example, several performance metrics have been proposed for robot swarms (Hamann, 2013; Valentini et al, 2015) and self-organizing systems (Kaddoum et al, 2010; Eberhardinger et al, 2017, 2018), but these methods are particular to decentralized systems and would not apply to centralized ones.
@@
In this paper, we study the question of comparative performance assessment across different types of multi-robot control structures, using sweep coverage as an example task. We test a series of approaches spanning from more centralized to more decentralized and compare their results when tested in the same environment under the same task conditions. We define the degree of (de)centralization of a multi-robot system according to the structure of communication and style of coordination among the robots, as well as the degree to which they use global information about the task, the environment, or each other.
@@ 2 Sweep coverage: related work
In static area coverage, robots occupy predominantly stationary positions distributed throughout the environment. Collectively, they monitor the entire environment at once. Although quite effective, static coverage can require too many robots when the environment is large. In sweep coverage, by contrast, robots are mobile. Instead of occupying stationary positions, robots move systematically through the environment and periodically visit all portions of it.
@@ 2.1.1 Centralized approaches
When organized in a centralized way, the environment can be decomposed into zones which robots are assigned to sweep individually (e.g., Rekleitis et al, 2008; Scherer et al, 2015). If the environment is decomposed a priori and path planning is completed offline, various forms of optimization can be used (Thabit and Mohades, 2018; Nazarahari et al, 2019; Yu and LaValle, 2016), often increasing efficiency but decreasing adaptability. If decomposition and planning are instead completed online, robots can build a shared reference map of the environment either before coverage begins (Mirzaei et al, 2011) or during coverage using broadcasting (Miki et al, 2018; Ge and Fua, 2005). Shared reference maps are also used without strict decomposition into zones, with maps updated by single-broadcast (Marjovi et al, 2009) or somewhat less efficiently by multi-broadcast (Albani et al, 2017). Approaches updating reference maps in these ways are often hybrid to some degree, rather than strictly centralized, as they include some aspects of decentralization.
@@
Figure 1: Illustration of obstacles in the environment. 2D illustrations of the environment, showing the grid cells used for analysis and the number of obstacles: (a) low difficulty, 100 obstacles; (b) medium difficulty, 200 obstacles; (c) high difficulty, 300 obstacles. The obstacles, arena, and grid cells are shown at their correct relative sizes.
@@ 3 Study design
The experiment arena used is a $4\times 4$ m2 enclosed square, decomposed into $16\times 16$ grid cells. For complete coverage, each of the 256 grid cells must be visited by the ground robots. The difficulty of this task is increased by adding small $4\times 4\times 2$ cm3 obstacles to the environment (see Fig. 1). In each trial, the obstacle positions and orientations are randomly selected with uniform distribution. We test arenas with the following three obstacle difficulties: 100 obstacles for low difficulty (i.e., 1% of the environment surface occupied), 200 obstacles for medium difficulty, and 300 obstacles for high difficulty.
@@
Figure 2: The five control approaches. Random walk: Ground robots explore the environment independently via randomized motion control, without any UAV supervisors. Decentralized with beacons: Ground robots explore the environment independently via randomized motion control, except when UAV beacons instruct them to vacate an area. Hybrid formation: The UAV “brain” conducts a reactive boustrophedon sweep and all UAVs give motion instructions to ground robots in their limited fields of view, using a self-organized hierarchical communication network (using SoNS). Centralized formation: One UAV with unlimited field of view conducts a reactive boustrophedon sweep and gives motion instructions to all ground robots using a predefined centralized communication network. Predetermined: After the environment is decomposed and paths are planned offline, one UAV gives ground robots predetermined motion instructions, using a predetermined centralized communication network.
@@ 4.2 Decentralized with beacons
4.3 Hybrid formation based on self-organizing nervous systems (SoNSs)
We use hybrid formation control based on self-organizing nervous systems (SoNS) Zhu et al (2024). Using SoNS, a heterogeneous swarm can self-organize into a dynamic ad-hoc communication network with a hierarchical structure in the form of a tree (the root of the tree is called brain, and the nodes directly connected to the brain are its children; for full details, see Zhu et al (2024)). Using this hierarchy, robots report sensing events and cede part of their autonomy to the brain robot (see Fig. 2). The brain robot determines its own motion trajectory and, as it moves, it acts as a motion reference for its children (i.e., the brain acts as a reference coordinate frame used to calculate the motion instructions sent to its children, which in turn act as motion references for their children). The robots in a SoNS can self-reconfigure into a new communication network and control hierarchy on the fly, for instance, in case of task change or brain failure (Zhu et al, 2024).
Algorithm 1 Random walk: ground robot.
@@ 4.3 Hybrid formation based on self-organizing nervous systems (SoNSs)
Algorithm 2 Decentralized with beacons: ground robot.
@@
Algorithm 3 Hybrid formation (with SoNS) and centralized formation methods: ground robot.
@@
11:else follow motion instructions received from a UAV
@@
Algorithm 4 Predetermined: ground robot. (See Alg. 5 in the Appendix for more details.)
@@
5:else follow motion instructions received from the UAV
@@
In our hybrid formation setup, there are some targets guiding the self-organization of the SoNS. The target communication network topology is a caterpillar tree—i.e., a tree in which all inner nodes are on one central path, to which each leaf node is connected. UAVs try to recruit ground robots and assign them to leaf nodes in the communication network. The ground robots complete the coverage task using instructions they receive from UAV supervisors and perform independent obstacle avoidance. The pseudocode for the ground robots is given in Algorithm 3; note that if ground robots temporarily lose connection with their UAV supervisors, they revert to the random walk behavior. UAV supervisors try to recruit each other, such that two of them become inner nodes and one becomes the brain node. The UAVs act as motion references for their children and direct their ground robot children into a line formation, as shown in Fig. 2, and prevent collisions between ground robots. The UAV that becomes the brain (see Zhu et al, 2024) reactively sweeps the environment using back-and-forth boustrophedon motions (Choset and Pignon, 1998), in the same manner as in (Jamshidpey et al, 2023). When the brain detects an arena boundary in front of it and shifts to the left, its motion calculation is based on its knowledge of the formation’s dimensions. Thus, the ground robots’ coverage of the environment is maximized regardless of how many ground robots are in the formation. A behavior flowchart for all UAVs in the SoNS hybrid formation is given in Fig. 3(a).
@@
(a) Hybrid formation: brain and non-brain UAVs.
(b) Centralized formation: UAV supervisor.
(c) Predetermined: UAV supervisor.
Figure 3: Sweep motion routines of the UAVs. Flowcharts for UAV sweep behaviors in the SoNS hybrid formation, centralized formation, and predetermined methods.
@@ 4.6 Experiment setup
Fiducial markers encoding unique IDs sit atop the ground robots and can be detected by the UAVs. The ground robots are incapable of localizing themselves, but the UAVs can track the relative positions and orientations of the ground robots in order to give them motion instructions. In the centralized formation and predetermined approaches, there is one supervisor UAV that can view the full arena and communicate with all ground robots. In the SoNS hybrid formation approach, a UAV can communicate with a ground robot if the ground robot is in its field of view, and two UAVs can communicate if they can see the same ground robot (for details, see Zhu et al, 2020). The three UAVs fly at a constant altitude and each UAV can view ground robots in an approximately $1.65\times 1.85$ m2 rectangular ground area (for details, see Zhu et al, 2020).
At the start of each experiment, the $4\times 4\times 2$ cm3 obstacles are distributed randomly throughout the $4\times 4$ m2 arena, but with obstacles that are very close to each other getting clustered side-by-side (in clusters not exceeding $k_{2}$ in number), and while ensuring at least a $20$ cm buffer between obstacles or obstacle clusters on at least one axis of the environment coordinate system ($x$ or $y$) and at least a $18$ cm buffer between obstacles and the outer boundary. The buffers prevent the ground robots from becoming stuck.
@@
Figure 4: Performance: Coverage completeness over time. Shaded area or bars represent the minimum to maximum; line represents the mean; dotted line indicates that some of the trials have completed the sweep, so are no longer included in the data for that time step. Random walk (gray), decentralized with beacons (green), hybrid formation (yellow), centralized formation (purple), and predetermined (red) approaches.
@@ 5.1.1 Coverage completeness
At the end of a sweep, the predetermined, centralized formation, and hybrid formation approaches all achieved perfect or nearly perfect coverage completeness (99.9% to 100% completeness on average), with slightly more outlier variation occurring in the hybrid formation approach (see Fig. 5(a)). The predetermined approach, however, was faster than the centralized and hybrid formation approaches, most noticeably in the condition with 100 obstacles. As the number of obstacles increased, the time advantage of the predetermined approach decreased (taking approximately 17% less time on average with 100 obstacles, versus approximately 6% less time on average with 300 obstacles). The decentralized approaches are very similar in terms of average coverage completeness and completeness variability, with the decentralized approach with beacons showing slightly greater deviations. Both decentralized approaches achieve noticeably lower completeness than the predetermined and formation approaches and noticeably greater completeness variability, but they still perform quite well (91.5% to 92.3% completeness on average). However, they take approximately twice as much time to complete a sweep. For all approaches, the number of obstacles has no clear effect on the average coverage completeness achieved.
@@ 5.1.2 Coverage uniformity
We define coverage uniformity as a measure of the variability of cell visit time. The best performing coverage approach would be one in which 1) all cells are visited and 2) they are all visited for an equal amount of time. In this ideal case, there would be no variability in cell visit time, resulting in a uniformity value of 00.
@@
At the end of one sweep, the predetermined approach has noticeably more uniform coverage (higher $p$) than the centralized and hybrid formation approaches, which in turn have noticeably more uniform coverage than the two decentralized approaches, for all obstacle quantities (see Fig. 5(b)). All approaches have low variability of coverage uniformity between trials, with the formation approaches having the least. In all approaches, coverage uniformity worsens slightly as the number of obstacles increases, with the predetermined approach worsening at a faster rate than the other approaches.
(a) No failure.
(b) 2 failures (25% failure).
(c) 4 failures (50% failure).
(d) 6 failures (75% failure).
Figure 6: Fault tolerance: Coverage completeness over time under varying loss conditions. Shaded area or bars represent the minimum to maximum; line represents the mean; dotted line indicates that some of the trials have completed the sweep, so are no longer included in the data for that time step. Random walk (gray), decentralized with beacons (green), hybrid formation (yellow), centralized formation (purple), and predetermined (red) approaches.
(a) Fault tolerance: Coverage completeness.
(b) Fault tolerance: Coverage uniformity.

### v4

@@ Abstract
In swarm robotics, decentralized control is often proposed as a more scalable and fault-tolerant alternative to centralized control. However, centralized behaviors are often faster and more efficient than their decentralized counterparts. In any given application, the goals and constraints of the task being solved should guide the choice to use centralized control, decentralized control, or a combination of the two. Currently, the exact trade-offs that exist between centralization and decentralization are not well defined. In this paper, we compare the performance of centralization and decentralization in the example task of sweep coverage, across five different types of multi-robot control structures: random walk, decentralized with beacons, hybrid formation control using self-organizing hierarchy, centralized formation control, and predetermined. In all five approaches, the coverage task is completed by a group of ground robots. In each approach, except for the random walk, the ground robots are assisted by UAVs, acting as supervisors or beacons. We compare the approaches in terms of three performance metrics for which centralized approaches are expected to have an advantage—coverage completeness, coverage uniformity, and sweep completion time—and two metrics for which decentralized approaches are expected to have an advantage—scalability (4, 8, or 16 ground robots) and fault tolerance (0%, 25%, 50%, or 75% ground robot failure). As expected, the results showed that the more centralized approaches greatly outperformed the decentralized ones in terms of coverage completeness, coverage uniformity, and sweep completion time. The decentralized approaches were less affected by robot failures and had better performance gains when the number of robots increased, but unexpectedly, these advantages only made their performance comparable to that of the more centralized approaches, not better than. Finally, we discuss future work on investigating additional conditions (e.g., bottlenecks, supervisor failures, and more complex environments), and on combining the advantages of both centralization and decentralization into one system.
@@ 1 Introduction
When developing multi-robot systems, a fundamental design choice is whether to use centralization, decentralization, or some combination of the two. Ideally, one would be able to guarantee that the selected control approach is the top performer for the targeted objectives. However, current methods to assess performance do not span different control approaches, making direct comparisons difficult. For example, several performance metrics have been proposed for robot swarms (Hamann, 2013; Valentini et al, 2015) and self-organizing systems (Kaddoum et al, 2010; Eberhardinger et al, 2017, 2018), but these methods are particular to decentralized systems and would not apply to centralized ones.
@@
In this paper, we compare aspects of performance across different types of multi-robot control structures, using sweep coverage as an example task. We test a series of approaches spanning from more centralized to more decentralized and compare their results when tested in the same environment under the same task conditions. We define the degree of (de)centralization of a multi-robot system according to the structure of communication and style of coordination among the robots, as well as the degree to which the robots use global information about the task, the environment, or each other.
@@ 2 Sweep coverage: related work
In static area coverage, robots or sensors occupy predominantly stationary positions distributed throughout the environment (Wang, 2011). Collectively, they monitor the entire environment without moving. In sweep coverage, by contrast, robots are mobile (Galceran and Carreras, 2013). Instead of occupying stationary positions, robots move systematically through the environment and periodically visit all portions of it. In this paper, we study sweep coverage.
@@ 2.1.1 Centralized approaches
When organized in a centralized way, the environment can be decomposed into zones which robots are assigned to sweep individually (e.g., Rekleitis et al, 2008; Scherer et al, 2015). If the environment is decomposed a priori and path planning is completed offline, various forms of optimization can be used (Thabit and Mohades, 2018; Nazarahari et al, 2019; Yu and LaValle, 2016), often increasing efficiency but decreasing adaptability. If decomposition and planning are instead completed online, robots can build a shared reference map of the environment either before coverage begins (Mirzaei et al, 2011) or during coverage using broadcasting (Miki et al, 2018; Ge and Fua, 2005). Shared reference maps are also used without strict decomposition into zones, with maps updated by single-broadcast (Marjovi et al, 2009) or somewhat less efficiently by multi-broadcast (Albani et al, 2017). Approaches that update reference maps in these ways are often hybrid to some degree, rather than strictly centralized, as they include some aspects of decentralization.
@@
Figure 1: Illustration of obstacles in the environment. 2D illustrations of the environment, showing the grid cells used for analysis and the number of obstacles: (a) low difficulty, 100 obstacles; (b) medium difficulty, 200 obstacles; (c) high difficulty, 300 obstacles. The obstacles (shown in black) are all the same size, but are sometimes positioned in clusters (see Sec. 4.6 for more information). The grid cells (shown in light gray) are used only for post-hoc analysis of the experiment results; robots do not receive information about the grid cells and cannot detect them. The obstacles, arena, and grid cells are shown at their correct relative sizes.
@@ 3 Study design
The experiment arena used is a $4\times 4$ m2 enclosed square, decomposed into $16\times 16$ grid cells (grid shown in light gray in Fig. 1). For complete coverage, each of the 256 grid cells must be visited by the ground robots. The difficulty of this task is increased by adding small $4\times 4\times 2$ cm3 obstacles to the environment (shown as small black squares in Fig. 1). In each trial, the obstacle positions and orientations are randomly selected with uniform distribution. We test arenas with the following three obstacle difficulties: 100 obstacles for low difficulty (i.e., 1% of the environment surface occupied), 200 obstacles for medium difficulty, and 300 obstacles for high difficulty.
@@
Figure 2: The five control approaches. Random walk: Ground robots explore the environment independently via randomized motion control, without any UAV supervisors. Decentralized with beacons: Ground robots explore the environment independently via randomized motion control, except when UAV beacons instruct them to vacate an area. Hybrid formation: The UAV “brain” conducts a reactive boustrophedon sweep and all UAVs give motion instructions to ground robots in their limited fields of view, using a self-organized hierarchical communication network (using the SoNS). Centralized formation: One UAV with unlimited field of view conducts a reactive boustrophedon sweep and gives motion instructions to all ground robots using a predefined centralized communication network. Predetermined: After the environment is decomposed and paths are planned offline, one UAV gives ground robots predetermined motion instructions, using a predetermined centralized communication network.
@@ 4.2 Decentralized with beacons
4.3 Hybrid formation based on the self-organizing nervous system (SoNS)
We use hybrid formation control based on the self-organizing nervous system (SoNS) Zhu et al (2024). Using the SoNS, a heterogeneous swarm can self-organize into a dynamic ad-hoc communication network with a hierarchical structure in the form of a tree (the root of the tree is called brain, and the nodes directly connected to the brain are its children; for full details, see Zhu et al (2024)). Using this hierarchy, robots report sensing events and cede part of their autonomy to the brain robot (see Fig. 2). The brain robot determines its own motion trajectory and, as it moves, it acts as a motion reference for its children (i.e., the brain acts as a reference coordinate frame used to calculate the motion instructions sent to its children, which in turn act as motion references for their children). The robots in a SoNS can self-reconfigure into a new communication network and control hierarchy on the fly, for instance, in case of task change or brain failure (Zhu et al, 2024).
Algorithm 1 - Ground robot behavior: Random walk.
@@ 4.3 Hybrid formation based on the self-organizing nervous system (SoNS)
Algorithm 2 - Ground robot behavior: Decentralized with beacons.
@@
Algorithm 3 - Ground robot behavior: Hybrid formation and centralized formation methods. For the accompanying UAV behaviors, see flowcharts in Figs. 3(a) and 3(b).
@@
11:else follow motion instructions from UAV    (See Fig. 3)
@@
Algorithm 4 - Ground robot behavior: Predetermined. (See Alg. 5 in the Appendix for more details.) For the accompanying UAV behavior, see flowchart in Fig. 3(c).
@@
5:else follow motion instructions from UAV    (See Fig. 3)
@@
In our hybrid formation setup, the target communication network topology is a caterpillar tree—i.e., a tree in which all inner nodes are on one central path, to which each leaf node is connected. UAVs try to recruit ground robots and assign them to leaf nodes in the communication network. The ground robots complete the coverage task using instructions they receive from UAV supervisors and perform independent obstacle avoidance. The pseudocode for the ground robots is given in Algorithm 3; note that if ground robots temporarily lose connection with their UAV supervisors, they revert to the random walk behavior. UAV supervisors try to recruit each other, such that two of them become inner nodes and one becomes the brain node. The UAVs act as motion references for their children and direct their ground robot children into a line formation, as shown in Fig. 2, and prevent collisions between ground robots. The UAV that becomes the brain (see Zhu et al, 2024) reactively sweeps the environment using back-and-forth boustrophedon motions (Choset and Pignon, 1998), in the same manner as in (Jamshidpey et al, 2023). When the brain detects an arena boundary in front of it and shifts to the left, its motion calculation is based on its knowledge of the formation’s dimensions. Thus, the ground robots’ coverage of the environment is maximized regardless of how many ground robots are in the formation. A behavior flowchart for all UAVs in the SoNS hybrid formation is given in Fig. 3(a).
@@
(a) UAV behavior: Hybrid formation (brain and non-brain UAVs), with boustrophedon sweep behavior highlighted in blue. For the accompanying ground robot behavior, see Alg. 3.
(b) UAV behavior: Centralized formation, with boustrophedon sweep behavior highlighted in blue. For the accompanying ground robot behavior, see Alg. 3.
(c) UAV behavior: Predetermined. For the accompanying ground robot behavior, see Alg. 4.
Figure 3: UAV behaviors. Flowcharts for the UAV behaviors in the (a) hybrid formation, (b) centralized formation, and (c) predetermined methods. In the hybrid formation and centralized formation methods, the boustrophedon sweep behavior of the UAVs is the same (see portion of the flowchart highlighted in blue): in both cases, the boustrophedon sweep is accomplished simply by shifting left and reversing direction when a boundary is met. In both cases, the dimensions of the target formation are fixed, and the UAVs are aware of those dimensions, so they shift approximately one formation-length to the left.
@@ 4.6 Experiment setup
Fiducial markers encoding unique IDs sit atop the ground robots and can be detected by the UAVs. The ground robots are incapable of localizing themselves, but the UAVs can track the relative positions and orientations of the ground robots in order to give them motion instructions. In the centralized formation and predetermined approaches, there is one supervisor UAV that can view the full arena and communicate with all ground robots. In the SoNS hybrid formation approach, a UAV can communicate with a ground robot if the ground robot is in its field of view, and two UAVs can communicate if they can see the same ground robot (for details, see Zhu et al, 2020). The three UAVs fly at a constant altitude, and each UAV can view ground robots in an approximately $1.65\times 1.85$ m2 rectangular ground area (for details, see Zhu et al, 2020).
At the start of each experiment, the $4\times 4\times 2$ cm3 obstacles are distributed randomly throughout the $4\times 4$ m2 arena (shown as small black squares in Fig. 1), but with obstacles that are very close to each other getting clustered side-by-side (in clusters not exceeding $k_{2}$ in number), and while ensuring at least a $20$ cm buffer between obstacles or obstacle clusters on at least one axis of the environment coordinate system ($x$ or $y$) and at least a $18$ cm buffer between obstacles and the outer boundary. The buffers prevent the ground robots from becoming stuck.
@@
Figure 4: Performance: Coverage completeness over time. Shaded area or bars represent the minimum to maximum; lines represent the mean; dotted lines indicate that some of the trials have completed the sweep, so are no longer included in the data for that time step. Random walk (gray), decentralized with beacons (green), hybrid formation (yellow), centralized formation (purple), and predetermined (red) approaches.
@@ 5.1.1 Coverage completeness
At the end of a sweep, the predetermined, centralized formation, and hybrid formation approaches all achieved perfect or nearly perfect coverage completeness (99.9% to 100% completeness on average), with slightly more outlier variation occurring in the hybrid formation approach (see Fig. 5a). The predetermined approach, however, was faster than the centralized and hybrid formation approaches, most noticeably in the condition with 100 obstacles. As the number of obstacles increased, the time advantage of the predetermined approach decreased (taking approximately 17% less time on average with 100 obstacles, versus approximately 6% less time on average with 300 obstacles). The decentralized approaches are very similar in terms of average coverage completeness and completeness variability, with the decentralized approach with beacons showing slightly greater deviations. Both decentralized approaches achieve noticeably lower completeness than the predetermined and formation approaches and noticeably greater completeness variability, but they still perform quite well (91.5% to 92.3% completeness on average). However, they take approximately twice as much time to complete a sweep. For all approaches, the number of obstacles has no clear effect on the average coverage completeness achieved.
(a) No failure.
(b) 2 failures (25% failure).
(c) 4 failures (50% failure).
(d) 6 failures (75% failure).
Figure 6: Fault tolerance: Coverage completeness over time under varying loss conditions. Shaded area or bars represent the minimum to maximum; lines represent the mean; dotted lines indicate that some of the trials have completed the sweep, so are no longer included in the data for that time step. Random walk (gray), decentralized with beacons (green), hybrid formation (yellow), centralized formation (purple), and predetermined (red) approaches.
(a) Fault tolerance: Coverage completeness.
(b) Fault tolerance: Coverage uniformity.
(a) Scalability: Coverage completeness.
(b) Scalability: Coverage uniformity.
(c) Scalability: Sweep completion time.
@@ 5.1.2 Coverage uniformity
We define coverage uniformity as a measure of the variability of cell visit time. The best performing coverage approach would be one in which 1) all cells are visited and 2) they are all visited for an equal amount of time. In this ideal case, there would be no variability in cell visit time, resulting in a uniformity value of 0.
@@
At the end of one sweep, the predetermined approach has noticeably more uniform coverage (higher $p$) than the centralized and hybrid formation approaches, which in turn have noticeably more uniform coverage than the two decentralized approaches, for all obstacle quantities (see Fig. 5b). All approaches have low variability of coverage uniformity between trials, with the formation approaches having the least. In all approaches, coverage uniformity worsens slightly as the number of obstacles increases, with the predetermined approach worsening at a faster rate than the other approaches.